A man pulls a 10-kg sled 66 meters along an angled hill with a force of 53 N, which elevates the man 28 meters above the bottom of the hill. The man then hops on his sled and slides from rest to the bottom of the hill back along his 66 meter path, during which a 102 N frictional force acts upon his sled. How much work in Joules does the man do on the sled in pulling the sled up the hill?

A man pulls a 10-kg sled 51 meters along an angled hill with a force of 75 N, which elevates the man 31 meters above the bottom of the hill. The man then hops on his sled and slides from rest to the bottom of the hill back along his 51 meter path, during which a 295 N frictional force acts upon his sled. How much work in Joules does the man do pulling the sled up the hill?
Answer #1

workdone by man

W = F d = 53* 66 = 3498 J

=======

b)

again

w = F d = 75 * 51 = 3825 J

=====

Comment before rate in case any doubt, will reply for sure.. goodluck

> Part A is incorrect

Abby Barnett Sun, Nov 21, 2021 9:12 AM